OpenStudy (anonymous):

Solve this.... with steps and explain me the answer x=root(4+root(4-root(4+root(4-.....) solve this ... Refer comment

OpenStudy (anonymous):

\[\sqrt{4+\sqrt{4-\sqrt{4+\sqrt{4.......}}}}\]

ganeshie8 (ganeshie8):

square both sides

ganeshie8 (ganeshie8):

\[\rm x = \sqrt{4+\sqrt{4-\sqrt{4+\sqrt{4-\cdots}}}}\]

ganeshie8 (ganeshie8):

\[\rm x^2-4 = \sqrt{4-\sqrt{4+\sqrt{4-\cdots}}}\]

ganeshie8 (ganeshie8):

square again

ganeshie8 (ganeshie8):

\[\rm \left(x^2-4\right)^2 = 4 - \sqrt{4+\sqrt{4-\cdots}}\]

ganeshie8 (ganeshie8):

Notice that infinite nexted radical on right hand side is just x

OpenStudy (anonymous):

the problem what I am facing is solving the 4th degree equation!

ganeshie8 (ganeshie8):

\[\rm \left(x^2-4\right)^2 = 4 - x\] solve x such that \(x \ge 0\) and \(x^2 - 4 \ge 0\)

OpenStudy (anonymous):

i had already derives the equation till the step you have shown

ganeshie8 (ganeshie8):

solving 4th degree eqn is indeed a pain

OpenStudy (anonymous):

isnt there a shorter method to calculate manually

ganeshie8 (ganeshie8):

you can work it by hand by factorizing it into two quadratics, but i don't think it is easy

OpenStudy (anonymous):

chalo thanks..!